There are 3 fathers each have 2 kids each how will they all sit together if there is only 7 chairs?
Answer by checkley79(3341)

3 PAIRS OF 1 FATHER & 2 KIDS
3(1+2)=3*3=9 PEOPLE.
2 OF THE KIDS SIT ON THEIR FATHER'S LAP.